OpenStudy (anonymous):

Solve Algebraically x-3y+4z=7 2x+2y-3z=-3 x-3y+z=-2

HanAkoSolo (jamierox4ev3r):

what variable are we solving for?

OpenStudy (anonymous):

I believe all of them. All the instructions say was solve Algebraically so I think I have to find one variable at a time to ultimately find them all

HanAkoSolo (jamierox4ev3r):

interesting...the best way to solve for a variable is, of course, to isolate it.

HanAkoSolo (jamierox4ev3r):

wait...does this problem mention anything about solving a system of equations?

HanAkoSolo (jamierox4ev3r):

because, if you have to solve for each variable...we're going to have to use a multitude of techniques. For example, let me show you an example of what I mean by "isolating the variable."

HanAkoSolo (jamierox4ev3r):

x-3y+4z=7 ^ this is your first equation, is it not? since x is already by itself, this should be relatively easy to solve for! x-3y+4z=7 +3y +3y ------------ x+4z=7+3y -4z -4z x=7+3y-4z We get that x is equal to 7+3y-4z by isolating x in the first equation. Make sense so far?

HanAkoSolo (jamierox4ev3r):

@BinkyBoo you following me so far? :)

OpenStudy (anonymous):

Yes I think I am

HanAkoSolo (jamierox4ev3r):

alright. So do you think you can solve for any other variables for any other equations? Hint: you can isolate for something in the 3rd equation

OpenStudy (anonymous):

couldn't I isolate the x again?

OpenStudy (shadowlegendx):

x y z, the alg2 fun xD

HanAkoSolo (jamierox4ev3r):

you could. but you could also isolate for z. The more things you isolate for, the easier of a time we'll have substituting them in later and lol @ShadowLegendX indeed

OpenStudy (anonymous):

So then if i have one isolated for x and then one for z then I have to get one for y still right?

HanAkoSolo (jamierox4ev3r):

true..but y is not so easily isolated. this is where we have to start substituting

HanAkoSolo (jamierox4ev3r):

so we're going to multiply the third equation by 2, which gives us 2x + 4y - 6z = 2 ^ then, we add this to the first equation: -2x + 3y + z = 3 2x + 4y - 6z = 2 --------------------- 7y - 5z = 5

HanAkoSolo (jamierox4ev3r):

^^now that we have that narrowed down to two variables, we need to multiply the third equation by -3. What does that give us?

HanAkoSolo (jamierox4ev3r):

@BinkyBoo

HanAkoSolo (jamierox4ev3r):

Need a little more guidance, perhaps? lol don't leave me hanging here

OpenStudy (anonymous):

I'm sorry I'm trying to make sure I wasn't missing some other instructions because it seemed so vague it would give us 3x-9y+3z=-6

OpenStudy (anonymous):

I think I am understanding now.

HanAkoSolo (jamierox4ev3r):

right! You just keep substituting, finding common variables until you're able to solve for your first variable! Once you have that variable, the rest should be piece of cake.

OpenStudy (anonymous):

AHHHHH THANK YOU SO MUCH!!!!!! :D

HanAkoSolo (jamierox4ev3r):

no problem! <3

HanAkoSolo (jamierox4ev3r):

think of it as one gigantic puzzle ^_^ LOL been a while since I had a problem like this to solve! But once you crack the code, it should come quite easily

OpenStudy (lochana):

@Jamierox4ev3r well explained!

HanAkoSolo (jamierox4ev3r):

thank you :) In fact @BinkyBoo I can share with you an even quicker way of doing this! I just found it out, working by scratch

HanAkoSolo (jamierox4ev3r):

so if you multiply the first equation by -1, what do you get?

HanAkoSolo (jamierox4ev3r):

you get -x+3y-z You can then add this to the 1st equation, which gives you the following: x-3y+4z=7 -x+3y-z=2 ---------- 3x=9 and here, you can easily solve for x! Now that you have a value for x, you can plug it in to the rest of the equations to solve for x and y ^_^ I hope you find this method even easier

OpenStudy (anonymous):

Oh my gosh! how did I not see that at first haha thank you you broke this down perfectly!!!

HanAkoSolo (jamierox4ev3r):

no problem x'D sometimes, negatives can be our best friends. Especially when they help eliminate variables that we don't want